The following information relates to the only product sold by Harper Company.

Sales price per unit $ 45
Variable cost per unit 27
Fixed costs per year 262,000



a. Compute the contribution margin ratio and the dollar sales volume required to break even.

b. Assuming that the company sells 20,000 units during the current year, compute the margin of safety (in dollars).
Answer #1

Contribution margin=Sales-Variable cost

=(45-27)=18 per unit

Contribution margin ratio=Contribution margin/Sales

(18/45)=40%

Breakeven sales=Fixed cost/Contribution margin ratio

=(262,000/0.4)=$655,000

Total sales=(20,000*$45)=$900,000

Hence margin of safety=Total sales-Breakeven sales

=(900,000-655,000)=$245,000